OpenStudy (anonymous):

What is the sum of 2 1/4 and 6 2/5

OpenStudy (anonymous):

8 13/20?

OpenStudy (anonymous):

What is the sum of 2 5 and 1 4?

OpenStudy (anonymous):

is it 8 and 13/20??

OpenStudy (anonymous):

25+14 = 20+19 = 1+38=39

OpenStudy (anonymous):

correct?